OpenStudy (shadowfiend):

Accidental release 08/14/2012: Chains and Passports

OpenStudy (shadowfiend):

Mostly today's release was about backend infrastructure still. We got into a situation where we had messages piling up in an intermediary server and we ran out of hard drive space for our database. Our database wasn't happy about it, when our database isn't happy our web tier isn't happy, when our web tier isn't happy you're not happy, when you're not happy the world explodes, AKA we had to get everything running again. Fortunately and completely coincidentally, today's release included a fix that will prevent the message pileup from happening again. There's still one intermittent bug we're trying to chase down that's causing server issues (that one may have triggered at about 2:30AM Eastern time today), but we're looking into it.

OpenStudy (shadowfiend):

Silly me, I forgot. @farmdawgnation finally gave our equation editor a fraction button. BOOYAH!

hero (hero):

Nice work :D

OpenStudy (farmdawgnation):

\[\frac{ farmdawg }{ awesome }\]

OpenStudy (saifoo.khan):

\[\frac{ fraction | button }{ nice }\]

OpenStudy (shadowfiend):

Hm. I'm not sure you folks have realized what fractions mean :D

OpenStudy (ash2326):

Thanks guys for an awesome work again:D

OpenStudy (kainui):

\[\frac{ about }{ time }baby\]

OpenStudy (karatechopper):

\[\frac{ :O }{ kewlio!}\]

OpenStudy (anonymous):

Thanks for getting this site back up, GREAT JOB GUYS KEEP UP THE AWESOME WORK!!! :)

OpenStudy (anonymous):

\[\frac{ thanks }{ guys }\]

OpenStudy (turingtest):

fractions mean legible algebra problems :) thanks farm!

OpenStudy (vishweshshrimali5):

\[\cfrac{thanks}{everyone}\]